find three consecutive odd integers such that their sum decreased by the second equals 50.
Answer by ewatrrr(24785)


Hi,
Let x represent the first of three consecutive odd integers
the other two would then be (x + 2) and (x + 4)
x + (x + 2) + (x + 4) -(x+2) = 50
2x + 4 = 50
2x = 46
x = 23
23,25,27 would be the three consecutive odd integers
checkingout answer
23 + 27 = 50